# Dynamical arrest in active nematic turbulence

### v2

@@ Acknowledgments
Supplementary Figure 4: Skeleton detection and pseudo-defect classification. 4, Grayscale plot of the Frank free energy density, with black corresponding to $\left|\nabla\theta\right|^{2}=0$ and white corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 4, Periodic expansion of (4) to avoid false detection of nodes along the original square boundaries (white outline). 4, Local adaptive binarization of (4), followed by removal of small fragmented components not counted as walls. 4, Recursive thinning of (4), then a step of dilation to fill up tiny gaps within some walls, an additional thinning step to recover the skeleton, and a pruning step to discard open branches shorter than a small threshold. 4, Using the MorphologicalTransform tool we detect and highlight the skeleton nodes. 4, We select only those nodes confined to the original periodic square. 4, For the SkeletonEndPoints, we compute the vectors pointing outwards of their associated branch (black arrows) by finding the nearest neighbouring point on the wall skeleton. 4, The wall polarity associated with each SkeletonEndPoint is shown in red (green) if its scalar product with (4) is positive (negative). 4, Branchpoints (blue nodes), endpoints (red nodes) and startpoints (green nodes) are plotted with their orientation given by the local wall polarity.
@@
Supplementary Figure 5: Labyrinths detection in arrested wall networks. 5, Grayscale plot of the Frank free energy density, with white corresponding to $\left|\nabla\theta\right|^{2}=0$ and black corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 5, Periodic expansion of (5) to capture connected domains that meander through the periodic square boundaries. 5, Local adaptive binarization of (5), followed by a filling operator to discard isolated black fragments. 5, Finding the connected morphological domains in (5). 5, Cropping out the periodic square domain from (5). 5, Applying our algorithm for relabeling domains that are connected through the left-right and top-bottom boundaries.
[…]
Supplementary Figure 6: Verification of the numerical scheme. 6, Consecutive snapshots from a simulation initiated with a small bending modulation about the uniformly aligned state: $\theta(\bm{r},t=0)=\epsilon\cos(2\pi x)$, with $\epsilon=0.01$. Parameter values were set to $R=1$, $\nu=-1.1$, $A=1000$, and $S=1$ (extensile stress). Numerical parameters were set to $N=64$, $dt=0.01$, and $D=0$ (zero noise). Color indicates the Frank energy, $|\nabla\theta|^{2}$, and white lines trace the director $\bm{n}$. 6, We plot the numerical data corresponding to a horizontal slice, $\theta(x,y=0)$, for increasing times, starting from $t=0$ (lightest gray) up to $t=40\tau_{\text{a}}$ (black) in intervals of $2.5\tau_{\text{a}}$. It is shown that the angle profile relaxes on the predicted 1D steady state (dashed blue), obtained for the same parameter values as explained in our recent work37. 6, The maximum angle is plotted as function of time in our simulation (black). First, it is shown that the modulation grows exponentially in time with the growth rate $\Omega$ matching the linear stability dispersion relation 37 (cyan). The gray inset is a log-scale plot comparing the simulation results with this explicit prediction for small perturbations. Additionally, it is demonstrated that the maximal angle in the simulation relaxes at long times to the value predicted by the 1D steady state 37 (dashed blue).
Supplementary Movie Captions
@@
Supplementary Movie 1 Strong large-scale turbulence in the contractile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
@@
Supplementary Movie 2 Arrested turbulence in the extensile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
@@
Supplementary Movie 3 Active turbulence in the absence of flow alignment (tumbling regime). The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Extended Data Fig. 1a and 1b, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
@@
Supplementary Movie 4 Evolution from stripes to turbulence (contractile aligning vs extensile aligning). The evolution of the Frank free energy density, $|\nabla\theta|^{2}$, in the contractile aligning case ($S\nu=+1.1$, left) and the extensile aligning case ($S\nu=-1.1$, right). The wavelength of the initial stripes—with splay walls on the left and bend walls on the right—was chosen to match the selected wavelength in the fully-developed turbulent state. In both cases, the stripes coarsen and then experience a ‘zigzag’ instability that folds the walls. In the contractile case (left), folded walls tend to dissolve and reappear perpetually. In the extensile case (right, Fig. 2), walls are more robust and tend to branch. As growing branches avoid other walls, the system reaches a grid-locked state that fluctuates only slightly. Parameters were set to $R=1$, $\nu=-1.1$, and $A=19692$ (chosen so that the system size roughly equals 6 times the selected wavelength). The interval between movie frames is 50 computational time iterations, corresponding to a time interval of $1\tau_{\text{a}}$. The total duration of the Movie is $540\tau_{\text{a}}$.
@@
Supplementary Movie 5 Evolution of the wall-network skeleton and its nodes. Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints shown as green, blue and red triangular nodes, respectively (see Fig. 3). The detection of these structures is described in Methods C and Extended Data Fig. 5. This movie demonstrates the pseudo-topological transitions mediating the ageing of an arrested wall network. Parameter values are as in Fig. 5. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $1715\tau_{\text{a}}$.
Supplementary Note:
Topological pseudo-charge of domain wall nodes
In arrested and nearly arrested states, the domain walls are organized in a tree-like network. One can associate a planar, directed graph to this network, with three types of nodes: startpoints, endpoints, and branchpoints, as described in Figs. 3 to 3. In the long-time asymptotics, the startpoints are only residual, and the ageing tree is mostly formed by endpoints and branchpoints (Fig. 5, Supplementary Movie 5). While these nodes are not topological defects of the nematic field, one can assign to them a pseudo-charge which is conserved and contains useful information of topological nature, as discussed in this section. We emphasize that this pseudo-topology is an effective emergent property of the dynamics, that holds only for the appropriate parameter regimes and time asymptotics. Therefore, pseudo-defects are dynamically stabilized but are not topologically stable in the sense that they can be removed by a continuous perturbation of the nematic field.
$\ell_{\text{a}}^{2}\partial_{s}^{2}\theta=\frac{S(1+\nu\cos 2\theta)(\sin 2\theta-\sin 2\bar{\theta})}{4+R+R\nu^{2}+2R\nu\cos 2\theta},$ (S1)
where $\bar{\theta}$ is the orientation of the nematic director at the center of the domain wall. This equation supports sigmoidal angle profiles with an interface of thickness of the order of the active length $\ell_{a}$ that continuously connects two plateaux of uniform angle. As the interfaces separate regions of uniform angle, they are referred to as domain walls, also called kinks and antikinks in other contexts.
For $S\nu>-1$, the two plateaux values of the nematic angle are $\theta=\bar{\theta}\pm\frac{\pi}{2}$, so the total rotation of the nematic across the wall is $\Delta\theta=\pm\pi$. For $S\nu<-1$, the plateaux angles correspond to the so-called Leslie angle defined by $01+\nu\cos 2\theta_{L}=$ 0, which amounts to an angle rotation of $\Delta\theta=\pm(\pi-\arccos{|\nu|^{-1}})$ ) across the wall. We remark that the total angle variation across domain walls is independent of both the active length $\ell_{\text{a}}$ and the viscosity ratio $R$. For a domain wall centered at $s=s_{i}$, we define the smooth step functions $\Theta^{\pm}(s-s_{i})$ that connect the plateaux values via a total angle increment $\Delta\theta$, which is positive for $\Theta^{+}$ and negative for $\Theta^{-}$. The angle varies at a scale $\sim\ell_{\text{a}}$ about $s_{i}$. Similarly, the derivative with respect to the spatial coordinate $s$ defines two peaked functions $\dot{\Theta}^{\pm}(s-s_{i})$ such that, in the sharp-wall limit,
$\lim_{\ell_{\text{a}}\rightarrow 0}\dot{\Theta}^{\pm}(s-s_{i})=\Delta\theta\,\delta(s-s_{i})\equiv\pm\frac{Q(\nu)}{2\pi}\delta(s-s_{i}).$ (S2)
Here, $Q(\nu)$ is the elementary pseudo-charge function, defined as
[…]
which gives, for $S\nu<0$,
$Q(\nu)=\frac{1}{2}\left(1-\frac{1}{\pi}{\cal H}(|\nu|-1)\arccos{|\nu|^{-1}}\right)>0,$ (S4)
where ${\cal H}$ is the Heaviside step function. The elementary pseudo-charge function $Q(\nu)$ takes the value $1/2$ in the tumbling regime, and it spans the interval from $1/2$ to $1/4$ as $|\nu|$ is increased beyond $1$. The functional form of $Q(\nu)$ is plotted in Fig. S1.
[…]
Figure S1: $Q$ as a function of $|\nu|$. Functional form of the elementary charge $Q(\nu)$. Positive pseudo-defects have charge $+Q$, and negative ones $-Q$. For the tumbling regime $Q=1/2$ and for the aligning regime $Q$ decreases monotonically from 1/2 to 1/4.
Similarly, for an isolated branchpoint we define the pseudo-charge as the rotation of the director in regions excluding the walls along a closed loop enclosing the branchpoint (Fig. 3). In this case, we must subtract the fast variation across the three walls intersected by the integration path. Since there is always one wall pointing towards the interior and two pointing towards the exterior, the pseudo-charge for a branchpoint, in the sharp-wall limit, will be $-Q(\nu)$.
$Q_{\Gamma}=(m_{\text{in}}-m_{\text{out}})Q(\nu),$ (S8)
In the case of domain walls with a small but finite thickness compared to typical wall separation, the elementary pseudo-charge will slightly differ from the limiting value defined in Eq. S3. In that case, consecutive walls partially overlap so that the angle profile does not reach the asymptotic value of the plateau, and the magnitude of the angle variation across the walls, $\Delta\theta$, is slightly smaller. This is the case in the wall patterns studied in this work, so in the main manuscript we refer to the values taken by the pseudo-topological charge only as $\simeq Q(\nu)$. We remark, however, that the precise value of the elementary charge $Q(\nu)$ is not really informative of the relevant topological structure, that is, how walls are connected, but accounts for the detailed structure of the walls themselves. This may be important to determine their dynamical properties, such as their interactions or their rigidity to deformation. However, it is only the integer counts of positive and negative pseudo-defects (or the counts of inward and outward wall intersections), i.e., the multiples of the elementary charge, that are informative of the overall qualitative structure inside a given domain.
In the sharp-wall limit $\ell_{\text{a}}\rightarrow 0$, the idea of a pseudo-defect is reminiscent of Dirac’s construction of magnetic monopoles, based on attaching to them the so-called Dirac string, which carries a quantized magnetic flux48. In that way, the magnetic monopole can be made consistent with Maxwell’s equations. For instance, the total magnetic flux across a closed surface enclosing the magnetic monopole would be zero, as required by Gauss’ theorem for the absence of net magnetic charge, because the flux generated by the monopole is canceled by that carried out by the string, which also crosses the surface. Similarly, in our problem, the wall attached to the pseudo-defect provides a localized (fast) rotation of the director that cancels the (slow) rotation outside the wall to ensure the lack of true topological charge. In this way, the problem satisfies the local constraint of zero topological charge while mimicking with pseudo-defects the director field around actual topological defects in a large region of space around them.

### v3

@@ E.1 Skeleton detection and pseudo-defect classification
Figure 13: Skeleton detection and pseudo-defect classification. 13, Grayscale plot of the Frank free energy density, with black corresponding to $\left|\nabla\theta\right|^{2}=0$ and white corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 13, Periodic expansion of (13) to avoid false detection of nodes along the original square boundaries (white outline). 13, Local adaptive binarization of (13), followed by removal of small fragmented components not counted as walls. 13, Recursive thinning of (13), then a step of dilation to fill up tiny gaps within some walls, an additional thinning step to recover the skeleton, and a pruning step to discard open branches shorter than a small threshold. 13, Using the MorphologicalTransform tool we detect and highlight the skeleton nodes. 13, We select only those nodes confined to the original periodic square. 13, For the SkeletonEndPoints, we compute the vectors pointing outwards of their associated branch (black arrows) by finding the nearest neighbouring point on the wall skeleton. 13, The wall polarity associated with each SkeletonEndPoint is shown in red (green) if its scalar product with (13) is positive (negative). 13, Branchpoints (blue nodes), endpoints (red nodes) and startpoints (green nodes) are plotted with their orientation given by the local wall polarity.
@@
Figure 14: Labyrinths detection in arrested wall networks. 14, Grayscale plot of the Frank free energy density, with white corresponding to $\left|\nabla\theta\right|^{2}=0$ and black corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 14, Periodic expansion of (14) to capture connected domains that meander through the periodic square boundaries. 14, Local adaptive binarization of (14), followed by a filling operator to discard isolated black fragments. 14, Finding the connected morphological domains in (14). 14, Cropping out the periodic square domain from (14). 14, Applying our algorithm for relabeling domains that are connected through the left-right and top-bottom boundaries.
E.2 Labyrinths detection
The process of labyrinth detection in arrested wall networks is depicted in Fig. 14. Here, the frames of interest are those in which no startpoints have been detected. Similarly to our skeleton detection scheme, we apply a pipeline that uses the image processing tools in Mathematica (TM).
Initially, an inverted plot of $\left|\nabla\theta\right|^{2}$ is rendered (Fig. 14). The color scheme and plot range are selected to aid the binarization of the gaps between the walls. To assist in identifying connected domains that frequently pass through the square boundaries, this plot is periodically expanded using the ImageAssemble function (Fig. 14). The LocalAdaptiveBinarize function is then applied with a neighborhood pixel range set to 10, followed by the FillingTransform function to eliminate isolated black fragments not considered as walls (Fig. 14).
Next, we find the connected morphological domains in the binarized image using the MorphologicalComponents function (Fig. 14). These distictnly labeled domains are then cropped to retain only those within the original periodic square (Fig. 14).
To relabel domains that connect through the left-right and top-bottom boundaries, we employ a custom algorithm that ensures continuity across the periodic square (Fig. 14). This step completes the process of detecting the unicursal labyrinths which span the system size.
E.3 Topological defects detection
The detection of true defects is more straightforward than that of pseudo-defects, since the tensorial order parameter $\mathbf{Q}$ provides clear local signatures of their appearance. Specifically, we rely on the scalar order parameter $s=2\sqrt{Q_{11}^{2}+Q_{12}^{2}}$ and the so-called topological defect density:
$\rho_{\text{top}}=\frac{1}{\pi}\left(\partial_{x}Q_{11}\partial_{y}Q_{12}-\partial_{y}Q_{11}\partial_{x}Q_{12}\right).$
At a defect core, $s$ falls to low values, approaching zero in the continuum limit, while $\rho_{\text{top}}$ develops local extrema whose sign indicates the topological charge. We identified core regions using a thresholding approach: a vertex was marked as part of a +1/2+1/2 defect core if $s<s_{0}$ and $\rho_{\text{top}}>r_{0}\max(\rho_{\text{top}})$, and as part of a 1/2-1/2 defect core if $s<s_{0}$ and $\rho_{\text{top}}<r_{0}\min(\rho_{\text{top}})$, with $0<s_{0}<1$ and $0<r_{0}<1$ as threshold parameters.
For each frame, we generated two binarized images based on these criteria and tiled them across periodic boundaries to obtain an added margin. Using the ComponentMeasurements function, we extracted centroids of connected regions, corresponding to the locations of positive and negative defect cores. We then filtered these to retain only those centroids that lie within the original simulation domain. We explored several values of $s_{0}$ and $r_{0}$ to achieve the closest agreement between the counts of +1/2+1/2 and 1/2-1/2 defects, as required by topological charge conservation. The detected defect positions were overlaid on plots of the director field (computed as the eigenvector of $\mathbf{Q}$) to visually confirm the results.
APPENDIX F SUPPLEMENTAL MOVIE CAPTIONS
@@ E.3 Topological defects detection
Movie S1   Strong large-scale turbulence in the contractile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the movie is $2500\tau_{\text{a}}$.
@@
Movie S2   Arrested turbulence in the extensile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the movie is $2500\tau_{\text{a}}$.
@@
Movie S3   Active turbulence in the absence of flow alignment (tumbling regime). The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Fig. S135, showing snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the movie is $2500\tau_{\text{a}}$.
@@
Movie S4   Evolution from stripes to turbulence (contractile aligning vs extensile aligning). The evolution of the Frank free energy density, $|\nabla\theta|^{2}$, in the contractile aligning case ($S\nu=+1.1$, left) and the extensile aligning case ($S\nu=-1.1$, right). The wavelength of the initial stripes—with splay walls on the left and bend walls on the right—was chosen to match the selected wavelength in the fully-developed turbulent state. In both cases, the stripes coarsen and then experience a ‘zigzag’ instability that folds the walls. In the contractile case (left), folded walls tend to dissolve and reappear perpetually. In the extensile case (right, Fig. 2), walls are more robust and tend to branch. As growing branches avoid other walls, the system reaches a grid-locked state that fluctuates only slightly. Parameters were set to $R=1$, $\nu=-1.1$, and $A=19692$ (chosen so that the system size roughly equals 6 times the selected wavelength). The interval between movie frames is 50 computational time iterations, corresponding to a time interval of $1\tau_{\text{a}}$. The total duration of the movie is $540\tau_{\text{a}}$.
@@
Movie S5   Evolution of the wall-network skeleton and its nodes. Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints shown as green, blue and red triangular nodes, respectively (see Fig. 3). The detection of these structures is described in Appendix E and Fig. 13. This movie demonstrates the pseudo-topological transitions mediating the ageing of an arrested wall network. Parameter values are as in Fig. 5. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the movie is $1715\tau_{\text{a}}$.
Movie S6   Agreement between the constrained $\theta$ model and the unconstrained Q-tensor model at low defect core size. Evolution of the flow (top panels) and the elastic energy density (bottom panels) in the $\theta$ model, compared with the $Q$-tensor model at low values of $\epsilon$. The behavior of the latter converges to the constrained case as $\epsilon$ decreases relative to $\ell_{\text{a}}$. See also Figs. 7 and 11. Other parameters were set to $R=1$, $\nu=-1$, $S=1$ and $A=10^{4}$. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $0.5\tau_{\text{a}}$. The total duration of the movie is $220\tau_{\text{a}}$.
Movie S7   Transition from dynamical arrest to defect-laden turbulence. Evolution of the flow (top panels) and the elastic energy density (bottom panels) in the $Q$-tensor model, with $\epsilon$ varied across the defect proliferation threshold. See also Fig. 7. Other parameters were set to $R=1$, $\nu=-1$, $S=1$ and $A=10^{4}$. The interval between movie frames is 200 computational time iterations, corresponding to a time interval of $1\tau_{\text{a}}$. The total duration of the movie is $600\tau_{\text{a}}$.
Movie S8   Unraveling dynamical arrest via a core-size quench. The left panel shows the elastic energy density, and the right panel shows a line-integral-convolution plot of the director, overlaid with the scalar order parameter (black indicating $s\to 0$). We first present the evolution of an arrested state generated at a low defect core size ($\epsilon=0.2\ell_{\text{a}}$). We then pause to highlight the localized energy peaks associated with branchpoint pseudo-defects. Next, the simulation is continued with a larger defect core size, beyond the threshold for defect nucleation ($\epsilon=0.35\ell_{\text{a}}$), showing that defect pairs nucleate at suspected hotspots. Finally, we replay the last segment with the true defects explicitly traced: yellow points are +1/2+1/2 defects, while cyan points 1/2-1/2 defects. Other parameters were set to $R=1$, $\nu=-1$, $S=1$ and $A=10^{4}$.
Movie S9   Enlarged defect-free extensile system. Evolution of the elastic energy density from simulations of the constrained $\theta$ model (Section D.5) in the extensile, rod-aligning regime ($S\nu=-1.1$). For reference, the movie begins with activity $A=3.2*10^{5}$ on a $256\times 256$ grid, as in Fig. 1, where the system rapidly develops an arrested state with persistent labyrinths spanning the entire domain. It then shows a larger system with $A=5.12*10^{6}$ on a $1024\times 1024$ grid, where more extensive labyrinths form but are frequently fractured, preventing persistent system-spanning connectivity.
Movie S10   Enlarged defect-free contractile system. Evolution of the elastic energy density from simulations of the constrained $\theta$ model (Section D.5) in the contractile, rod-aligning regime ($S\nu=+1.1$). The movie begins with activity $A=3.2*10^{5}$ on a $256\times 256$ grid, as in Fig. 1. It then shows a larger system with $A=5.12*10^{6}$ on a $1024\times 1024$ grid.
APPENDIX G SUPPLEMENTAL ANIMATION CAPTIONS
Animation S1   Startpoint-endpoint pair birth and annihilation via wall inception and wall dissolution. This animation illustrates both the forward and inverse processes related to Fig. 3.
Animation S2   Endpoint-startpoint pair birth and annihilation via local wall deletion and wall completion. This animation illustrates both the forward and inverse processes related to Fig. 3.
Animation S3   Branchpoint-endpoint pair birth and annihilation via wall branching and wall retraction. This animation illustrates both the forward and inverse processes related to Fig. 3.
Animation S4   Branchpoint to startpoint transition by wall disjoining and startpoint to branchpoint transition by wall joining. This animation illustrates both the forward and inverse processes related to Fig. 3.
Animation S5   T1 transitions of connected branchpoints. This animation illustrates both the forward and inverse processes related to Fig. 3.